OpenStudy (anonymous):

find dy/dx of y=ln(2+(sqrt(x))

OpenStudy (turingtest):

chain rule: \[dy/dx=[1/(2+\sqrt{x})](1/2)x^{-1/2}=1/[2\sqrt{x}(2+\sqrt{x})]=1/(4\sqrt{x}+2x)\]